# Development and Characterization of MPGD-based Transition Radiation Detectors

### v1

@@ Abstract
Transition Radiation Detectors (TRDs) are useful for electron identification and hadron suppression in high energy nuclear and particle physics experiments. Conventional wire-chamber TRDs face operational limitations due to space charge effects, motivating the replacement of the amplification stage with MicroPattern Gaseous Detectors (MPGDs). In this work, we explore different MPGD technologies - Gas Electron Multiplier (GEM), Micro-Mesh Gaseous Structure (Micromegas), and Resistive Micro-Well (RWELL) - as alternative TRD amplification stages. We report on the design, construction, and in-beam characterization of three MPGD-based TRD prototypes exposed to 3–20 GeV mixed electron–hadron beams at the Fermilab Test Beam Facility (FTBF) and at the CERN SPS H8 beamline. Each detector consisted of a multi-layered radiator, an approximately 2 cm deep drift region, an MPGD amplification stage optimized for X-ray transition radiation detection in a 90%10% ratio of Xenon and CO2 (Xe:CO2 90:10), and a two-dimensional readout. The GEM-based TRD prototype achieved a pion suppression factor of about 8 at 90% electron efficiency at FTBF, while the Micromegas-based prototype - with an added GEM preamplification layer - demonstrated improved gain stability and clear TR photon discrimination at CERN. The RWELL prototype achieved stable operation but limited gain. Geant4-based studies confirmed the observed trends and highlighted the sensitivity of the TR yield to cathode material and radiator configuration. These studies represent the first in-beam measurements of Micromegas- and RWELL-based TRDs, along with discussion of the performance capabilities of a triple-GEM-TRD. The results demonstrate the feasibility of MPGDs as scalable, high-rate amplification structures for next-generation TRD applications.
@@ 1 Introduction
Transition Radiation Detectors (TRDs) are well suited for this task due to their ability to separate electrons from hadrons over a broad momentum range (1–150 GeV/) PDG. In this momentum range, only electrons produce X-ray transition radiation (TR) in appropriate radiators, suitable for detection in heavy gases in the 3–40 keV energy region. Due to its low material budget, TRD can be combined with other electron identification methods, such as calorimetry, for improved electron/hadron separation and cross-calibration of the two detector systems. Both regular foil radiators and irregular fleece radiators are commonly used, each offering tradeoffs in photon yield, spectral distribution, and material budget.
A typical gaseous TRD consists of a radiator, gas volume, and amplification stage to detect the ionization deposited in the gas. Traditional TRD wire-based amplification structures have experienced performance issues due to space charge effects limiting their performance in HENP experiments TRD. This motivates the replacement of the amplification stage with MicroPattern Gaseous Detector (MPGD) technologies such as Gas Electron Multiplier (GEM) GEM, Micro-Mesh Gaseous Structure (Micromegas) MMG, and Resistive Micro-Well (RWELL) URW. These technologies offer excellent rate capability, reduced ion backflow, improved energy resolution, and simplified scalability to large systems.
Previous efforts at Thomas Jefferson National Accelerator Facility (Jefferson Lab) have demonstrated a proof-of-principle GEM-based TRD prototype, achieving an electron–pion (e/) rejection factor greater than 5 at 90% electron efficiency, along with three-dimensional charged-particle tracking capability NIMGEMTRD1. In this paper, we build on this previous work to explore Micromegas and RWELL as amplification choices in a TRD application for the first time. Both MPGD options possess practical advantages such as lower operating voltages, fewer high-voltage channels, more uniform gain over larger surface area, and simplified construction compared to GEMs, while retaining comparable energy resolution.
We present the design, development, and beam-test characterization of several MPGD-based TRD prototypes. In Section 2, we introduce the concept of a triple-GEM-, RWELL-, and Micromegas- based TRD and describe the design and construction of prototypes. Section 3 describes the in-beam studies of these prototypes in mixed hadron-electron beams at the Fermilab Test Beam Facility (FTBF) and the CERN SPS H8 beamline. Section 4 presents findings on gain stability, detector efficiency, timing performance, and e/ separation while Section 5 discusses the implications of these results and complementary Geant4-based simulations. Section 6 summarizes our conclusions on these different MPGD technologies as amplification layers in a TRD concept and outlines future tests planned for these novel detectors.
@@ 2 Prototype Development and Design
Optimizing TRD performance requires careful consideration of several detector parameters: the cathode material and thickness to maximize transmission of soft X-rays without compromising stability of operation, the drift region length to absorb higher-energy TR photons, and the radiator design to maximize TR production while minimizing self-absorption. The design concept for the TRD prototypes developed in this study focuses on simultaneously maximizing the probability of detecting TR photons and minimizing material budget and dead area. Each detector consists of a radiator upstream of a gaseous drift region followed by an MPGD amplification stage and a two-dimensional strip readout plane. The drift region, ranging from 20–28 mm in length, was chosen to efficiently absorb TR photons in the 3–40 keV range while maintaining stable electric field operation. Cathode foils were selected to minimize soft X-ray absorption without degrading mechanical and electrical robustness. Both fleece-type and foil-type radiators were tested, enabling a direct comparison of irregular versus regular radiator geometries for TR generation. The signal amplification stage was realized with the following MPGD technologies: GEMs, Micromegas, and RWELLs.
@@ 2.1 Triple-GEM-TRD
The GEM-TRD serves as the reference technology for these tests, previously studied in detail at Jefferson Lab NIMGEMTRD1. The detector employs a standard x cm2 triple-GEM stack with a drift gap of 21 mm, and transfer and induction gaps of 2 mm. After initial beam-tests at FTBF, the design was modified by replacing the original 25 m Kapton entrance window and Chromium-clad-Kapton cathode with a 5 m Cu foil on 55 m Kapton, eliminating a dead gas gap that had led to absorption of soft TR photons. The initial prototype design is referred to as GEM-TRD_v1, while the modified design with the copper-clad-Kapton cathode is referred to as GEM-TRD_v2; schematics of these two constructions are shown in Figure 1. For both designs of the GEM-TRD, a voltage divider was used for the triple-GEM amplification structure as a means of reducing the necessary number of HV channels from seven to two. Nominal voltage settings and electric field values are provided in A.
@@ 2.2 Micromegas-Based TRD
A Micromegas-TRD prototype was constructed at Vanderbilt University and tested in-beam at FTBF. The readout board employed a two-dimensional orthogonal X–Y strip layout with 256 strips per dimension at 400 m pitch, and the drift region was 28 mm. The cathode was a fine stainless steel mesh of 18 m thickness mounted approximately 1 mm below a 25 m Mylar entrance window. Following the FTBF tests, the prototype design was modified by introducing a single GEM foil upstream of the Micromegas mesh, separated by a 2.5 mm transfer gap and thereby reducing the drift gap to 25.5 mm. The GEM preamplification stage was introduced to improve the overall gain stability and reduce the probability of discharges, which caused limitations in stable operation during the FTBF tests. The introduction of a GEM preamplification layer also lead to the use of a voltage divider for the amplification structure to reduce the number of input HV channels to two. Nominal voltage settings and electric field values are given in B. The Micromegas+GEM-TRD also utilized an entrance window of a 5 m Cu foil on 55 m Kapton, for direct comparability to the GEM-TRD_v2 design and in order to remove the non-sensitive gas gap present in the Micromegas-TRD design. Figure 2 displays a cross section of both Micromegas-based prototype constructions.
@@
A RWELL-TRD prototype was assembled and tested at Jefferson Lab prior to in-beam tests. Its distinguishing feature is the use of a chromium-coated polyimide cathode foil (200 nm Cr on 50 m Kapton), reducing the material budget seen by TR photons while maintaining operational robustness. The drift gap was 25 mm. The amplification stage consisted of a resistive Micro-Well structure coupled to an X–Y strip layout with 128 strips per dimension, equating to a 800 m strip pitch as described in capaSh_urwell2022. Figure 3 illustrates the detector construction.
@@ 2.4 Readout Electronics and DAQ System
Fast signal collection is essential for TRDs in order to separate clusters along the track and distinguish the TR photon-induced signals from energy deposited in the gaseous drift region by all charge particles. For these prototypes, flash ADC modules (fADC-125) fadc with 8 ns sampling time developed by Jefferson Lab were used, and signals were amplified with fast (10 ns peaking time) GAS-II preamplifiers GAS2 before digitization. These fast front-end electronics require higher detector gas gain than traditional APV25-based readout apv25, but are critical for characterizing the timing and amplitude distributions of TR photon clusters.
3 In Beam Tests
The three detectors described in the previous section were tested in mixed hadron-electron beam at FTBF, where neither the Micromegas- nor the RWELL-TRD reached sufficient gain to enable a decisive measurement of hadron suppression. This limited gain stability observed at FTBF motivated the modification of the Micromegas prototype, where a GEM preamplification stage was implemented. Both the triple-GEM and Micromegas-based prototypes had their entrance window and cathode modified to be combined into one cohesive layer following the FTBF campaign. This alteration was done as a way to eliminate excess absorption of soft TR photons in the non-sensitive gas gap between these layers.
@@ 3.1 Data Selection
For in-beam measurements, several observables are recorded using fADC-125 electronics: pulse amplitude (corresponding to deposited energy) and drift time of clusters across the drift gap. Event-by-event analyses were performed to study the detector response for electrons and hadrons, enabling determination of e/ separation power. Figure 4 shows an example of raw waveforms collected for both the GEM-TRD and the Micromegas+GEM-TRD. For analysis of in-beam measurements, tight selections are applied to external particle identification (PID) detectors to assure high sample purity of charged particle types.
3.2 FTBF Experimental Setup
In May 2023, the GEM-TRD_v1, Micromegas-TRD, and RWELL-TRD prototypes were tested at FTBF in mixed electron–hadron beams of 3 GeV and 10 GeV. Electron fractions were approximately 70% at 3 GeV and 50% at 10 GeV. A standard triple-GEM tracker with better than m spatial resolution was placed upstream of the setup in order to provide a track reference. It utilized APV25 front-end ASICs to shape the signals from all electrode channels, which were subsequently digitized using a Scalable Readout System (SRS) srs. Triggering was accomplished by two scintillators overlapped in front of the active area of the detectors. The TRD prototypes were operated in a 90%10% ratio of Xenon and CO2 (Xe:CO2 90:10). Cherenkov detectors in the beamline were tuned to discriminate electrons from hadrons and used for external electron and pion sample selection. Additional PID was provided by a 7-cell lead-tungsten electromagnetic calorimeter supplied by Jefferson Lab. Fig. 5 shows the experimental setup.
Two radiator types were used in order to perform tests comparing the TR-yield of different materials at FTBF: fleece radiator material with irregular structure from the HERMES and ZEUS experiments hermes zeus and radiators constructed with regularly spaced Mylar foils having 12.5 m thickness separated by netted spacers of 300 m thickness. The fleece material had a density of roughly 0.09 g/cm3, and the Mylar foil radiators had a density of approximately 0.065 g/cm3.
3.3 CERN SPS Experimental Setup
In July 2024, tests of the GEM-TRD_v2 and Micromegas+GEM-TRD were carried out at CERN SPS H8 beam line with 20 GeV mixed electron–hadron beams. The setup included three standard triple-GEMs for precise tracking, an upstream Cherenkov detector for PID, upstream scintillators for timing and triggering, and a downstream single-cell Pb-glass calorimeter for additional PID. Figure 6 illustrates the CERN setup. Multiple radiator configurations were tested; for fleece material, both 15 cm and 23 cm lengths were used for the GEM-TRD_v2 and the Micromegas+GEM-TRD. Regarding Mylar foil, 15 cm of 30 m-thick foil with 240 m spacing was used on the GEM-TRD, while for the Micromegas+GEM-TRD 10 cm of 30 m-thick foil with 200 m spacing combined with 10 cm of 25 m-thick foil with 200 m spacing for a total length of 21 cm was used.

### v2

@@ Abstract
Transition Radiation Detectors (TRDs) are useful for electron identification and hadron suppression in high energy nuclear and particle physics experiments. Conventional wire-chamber TRDs face operational limitations due to space charge effects, motivating the replacement of the amplification stage with MicroPattern Gaseous Detectors (MPGDs). This work explores different MPGD technologies – Gas Electron Multiplier (GEM), Micro-Mesh Gaseous Structure (Micromegas), and Resistive Micro-Well (RWELL) – as alternative TRD amplification stages. We report on the design, construction, and in-beam characterization of multiple MPGD-based TRD prototypes exposed to 3–20 GeV mixed electron–hadron beams at the Fermilab Test Beam Facility and at the CERN SPS H8 beamline. Each detector consisted of a multi-layered radiator, an approximately 2 cm deep drift region, an MPGD amplification stage optimized for X-ray transition radiation detection in a Xe:CO2 (90:10) gas mixture, and a two-dimensional readout. The GEM-based TRD prototype achieved a pion suppression factor of about 8 at 90% electron efficiency, while the Micromegas-based prototype – with an added GEM preamplification layer – demonstrated improved operational stability and clear TR photon discrimination. The RWELL prototype achieved stable operation but limited signal gain.Geant4-based studies confirmed the observed trends and highlighted the sensitivity of the TR yield to cathode material and radiator configuration. These studies represent the first in-beam measurements of Micromegas- and RWELL-based TRDs, along with discussion of the performance capabilities of a triple-GEM-TRD. The results demonstrate the feasibility of MPGDs as scalable, high-rate amplification structures for next-generation TRD applications.
@@ 1 Introduction
Transition Radiation Detectors (TRDs) are well suited for this task due to their ability to separate electrons from hadrons over a broad momentum range (1–150 GeV/) [15]. In this momentum range, only electrons produce X-ray transition radiation (TR) in appropriate radiators, suitable for detection in heavy gases in the 3–40 keV energy region. Due to its low material budget, TRD can be combined with other electron identification methods, such as calorimetry, for improved electron/hadron separation and cross-calibration of the two detector systems. Both regular foil radiators and irregular fleece radiators are commonly used, each offering tradeoffs in photon yield, spectral distribution, and material budget.
A typical gaseous TRD consists of a radiator, gas volume, and amplification stage to detect the ionization deposited in the gas. The ALICE TRD [9] is one example of a modern TRD with typical drift chamber amplification. Space charge effects in traditional wire-based amplification structures have been previously reported [3] [16]. While able to be mitigated, these effects have resulted in performance issues for TRDs that have limited their application in HENP experiments. This motivates the replacement of the amplification stage with MicroPattern Gaseous Detector (MPGD) technologies such as Gas Electron Multiplier (GEM) [17], Micro-Mesh Gaseous Structure (Micromegas) [11], and Resistive Micro-Well (RWELL) [6]. These technologies offer excellent rate capability, reduced ion backflow, improved energy resolution, and simplified scalability to large systems.
Previous efforts at Thomas Jefferson National Accelerator Facility (Jefferson Lab) have demonstrated a proof-of-principle GEM-based TRD prototype, achieving an electron–pion (e/) rejection factor greater than 5 at 90% electron efficiency, along with three-dimensional charged-particle tracking capability [4]. In this paper, we build on this previous work to explore Micromegas and RWELL as amplification choices in a TRD application for the first time. Both MPGD options possess practical advantages such as lower operating voltages, fewer high-voltage channels, more uniform gain over larger surface area, and simplified construction compared to GEMs, while retaining comparable energy resolution.
We present the design, development, and beam-test characterization of several MPGD-based TRD prototypes. In Section 2, we introduce the concept of triple-GEM-, RWELL-, and Micromegas-based TRDs and describe the design and construction of prototypes. Section 3 describes the in-beam studies of these prototypes in mixed hadron–electron beams at the Fermilab Test Beam Facility (FTBF) and the CERN SPS H8 beamline. Section 4 presents findings on operational stability, detector efficiency, timing performance, and e/ separation while Section 5 discusses the implications of these results and complementary Geant4-based simulations. Section 6 summarizes our conclusions on these different MPGD technologies as amplification layers in a TRD concept and outlines future tests planned for these novel detectors.
@@ 2 Prototype Development and Design
Optimizing TRD performance requires careful consideration of several detector parameters: the entrance window material and thickness to maximize transmission of soft X-rays without compromising stability of operation, the drift region length to absorb higher-energy TR photons, and the radiator design to maximize TR production while minimizing self-absorption. The design concept for the TRD prototypes developed in this study was guided by the goal of simultaneously maximizing the probability of detecting TR photons while minimizing material budget and dead area, within the practical constraints of iterative R&D and prototyping. Each detector consists of a radiator upstream of a gaseous drift region followed by an MPGD amplification stage and a two-dimensional strip readout plane. The drift region, ranging from 20–28 mm in depth, was chosen to efficiently absorb TR photons in the 3–40 keV range while maintaining stable electric field operation; previous MC simulations showed that expansion beyond this range does not significantly improve pion suppression performance [4]. Cathode foils were selected with consideration for balancing soft X-ray absorption without compromising overall mechanical and electrical robustness. Both fleece-type and foil-type radiators were tested, enabling a direct comparison of irregular versus regular radiator geometries for TR generation. The signal amplification stage was realized with the following MPGD technologies: GEMs, Micromegas, and RWELLs. Table 1 displays a comparison of major design parameters between each of the tested prototypes.
| Parameter | GEM–TRD | GEM–TRD | Micromegas | Micromegas | RWell |
| _v1 | _v2 | –TRD | +GEM–TRD | –TRD | |
| Window material | Kapton | Kapton | Mylar | Kapton | Kapton |
| Window thickness [m] | 25 | 55 | 25 | 55 | 50 |
| Dead gas gap [m] | 400 | 0 | 1000 | 0 | 0 |
| Cathode material | Cr | Cu | Stainless Steel | Cu | Cr |
| Cathode thickness [m] | 0.2 | 5 | 18 | 5 | 0.2 |
| Drift gap depth [mm] | 21 | 19.9 | 28 | 25.5 | 25 |
| Chamber depth [mm] | 28 | 26.5 | 29.5 | 28.5 | 25.5 |
@@ 2.1 Triple-GEM-TRD
The GEM-TRD serves as the reference technology for these tests and was previously studied in detail at Jefferson Lab [4]. The detector employs a standard cm2 triple-GEM amplification stack with transfer and induction gaps of 2 mm. Charge readout is provided by a two-dimensional orthogonal X–Y strip layout comprising of 256 strips per dimension with a pitch of 400 m. Two separate GEM-TRD prototype designs were evaluated. The first, referred to as GEM-TRD_v1, retained the configuration described in [4] originally tested at Jefferson Lab. GEM-TRD_v1 was tested to extend the original proof-of-principle studies to hadron beams at FTBF. While GEM-TRD_v1 successfully demonstrated TRD operation, its design incorporates a 400 m dead gas gap immediately following the entrance window. If scaled to larger detector areas, this dead region is expected to increase in depth due to entrance-window bulging under operating gas pressure, leading to enhanced absorption of soft transition radiation photons in the heavy gas before reaching the active drift region. To isolate and eliminate this effect, a second triple-GEM prototype, referred to as GEM-TRD_v2, was constructed without an entrance-window-induced dead gas gap. In this design, the 25 m Kapton entrance window present in GEM-TRD_v1 was removed entirely and a 5 m copper cathode foil supported on 55 m Kapton was installed. The choice of 5 m Cu reflects a more conventional choice for cathode material and avoids potential mechanical and electrical stability concerns associated with the ultra-thin 200 nm Cr cathode employed in the GEM-TRD_v1 design. Schematics of these two constructions are shown in Figure 1. For both designs of the GEM-TRD, a voltage divider was used for the triple-GEM amplification structure as a means of reducing the necessary number of HV channels from seven to two. Nominal voltage settings and electric field values are provided in B.
@@ 2.2 Micromegas-Based TRD
A Micromegas-TRD prototype was constructed at Vanderbilt University and tested in-beam at FTBF. The readout board is of the same layout and design as the GEM-TRD, employing a two-dimensional orthogonal X–Y strip layout with 256 strips per dimension at 400 m pitch. The drift region was 28 mm, and the cathode was a fine stainless steel mesh of 18 m thickness mounted approximately 1 mm below a 25 m Mylar entrance window. Following the FTBF tests, the prototype design was modified by introducing a single GEM foil upstream of the Micromegas mesh, separated by a 2.5 mm transfer gap and thereby reducing the drift gap to 25.5 mm. The GEM preamplification stage was introduced to improve the overall signal gain and reduce the probability of discharges, which caused limitations in stable operation during the FTBF tests. The introduction of a GEM preamplification layer also lead to the use of a voltage divider for the amplification structure to reduce the number of input HV channels to two. Nominal voltage settings and electric field values are given in B. The Micromegas+GEM-TRD also utilized an entrance window of a 5 m Cu foil on 55 m Kapton, for direct comparability to the GEM-TRD_v2 design and in order to remove the non-sensitive gas gap present in the Micromegas-TRD configuration. Figure 2 displays a cross section of both Micromegas-based prototype constructions.
@@
A RWELL-TRD prototype was assembled and tested at Jefferson Lab prior to in-beam tests. Its distinguishing feature is the use of a chromium-coated polyimide cathode foil (200 nm Cr on 50 m Kapton), reducing the material budget seen by TR photons while maintaining operational robustness. The drift gap was 25 mm. The amplification stage consisted of a resistive Micro-Well structure coupled to an X–Y strip layout with 128 strips per dimension, equating to a 800 m strip pitch as described in [12]. Figure 3 illustrates the detector construction.
@@ 2.4 Readout Electronics and DAQ System
Fast signal collection is essential for TRDs in order to separate clusters along the track and distinguish the TR photon-induced signals from energy deposited in the gaseous drift region by all charged particles. For these prototypes, flash ADC modules (fADC-125) [19] with 8 ns sampling time developed by Jefferson Lab were used, and signals were amplified with fast (10 ns peaking time) GAS-II preamplifiers [5] before digitization. These fast front-end electronics require higher detector gas gain than commonly used APV25-based readout [10], but are critical for characterizing the timing and amplitude distributions of TR photon clusters. Regarding the readout electronics setup, 240 out of 256 strips on the X-plane of each detector are connected individually to ADC modules and read out in parallel, while all 256 strips of the Y-plane are read out serially with the SRS.
3 In-Beam Tests
The various detectors described in the previous section were tested in mixed hadron–electron beams at FTBF and CERN SPS. Neither the Micromegas-TRD nor the RWELL-TRD could reach sufficient signal gain to enable a decisive measurement of hadron suppression. This observed signal gain limitation of a single amplification layer motivated the modification of the Micromegas prototype into a hybrid MPGD design, where a GEM preamplification stage was implemented. The Micromegas+GEM-TRD also utilized an entrance window of a 5 m Cu foil adhered to 55 m Kapton, for direct comparability to the GEM-TRD_v2 design as discussed in Section 2.
3.1 FTBF Experimental Setup
In May 2023, the GEM-TRD_v1, Micromegas-TRD, and RWELL-TRD prototypes were tested at FTBF in mixed electron–hadron beams of 3 GeV and 10 GeV. Electron fractions were approximately 70% at 3 GeV and 50% at 10 GeV. A standard triple-GEM tracker with better than m spatial resolution was placed upstream of the setup in order to provide a track reference. It utilized APV25 front-end ASICs to shape the signals from all electrode channels, which were subsequently digitized using a Scalable Readout System (SRS) [14]. Triggering was accomplished by two scintillators overlapped in front of the active area of the detectors. The TRD prototypes were operated in a Xe:CO2 (90:10) gas mixture. Cherenkov detectors in the beamline were tuned to discriminate electrons from hadrons and used for external electron and pion sample selection. Additional particle identification (PID) was provided by a 7-cell lead–tungsten electromagnetic calorimeter supplied by Jefferson Lab. Figure 5 shows the experimental setup. The calorimeter provides a pion suppression factor of roughly 100, leading to sample purity of about 99%; the Cherenkov detectors provide an additional redundancy in external PID, resulting in adequate sample purity.
Two radiator types were used in order to perform tests comparing the TR-yield of different materials at FTBF: fleece radiator material with irregular structure from the HERMES and ZEUS experiments [1] [18] and radiators constructed with regularly spaced Mylar foils having 12.5 m thickness separated by netted spacers of 300 m thickness. The fleece material had a density of roughly 0.09 g/cm3, and the Mylar foil radiators had a density of approximately 0.065 g/cm3. Table 2 describes the radiator parameters.
3.2 CERN SPS Experimental Setup
In July 2024, tests of the GEM-TRD_v2 and Micromegas+GEM-TRD were carried out at CERN SPS H8 beam line with 20 GeV mixed electron–hadron beams. The setup included three standard triple-GEMs for precise tracking, an upstream Cherenkov detector for PID, upstream scintillators for timing and triggering, and a downstream single-cell Pb-glass calorimeter for additional PID. Figure 6 illustrates the CERN setup. High purity trigger signals corresponding to charged particle species were selected with information from the calorimeter and used in the setup as trigger flags. Trigger signal contamination of the other species for each particle type was below [2]. Multiple radiator configurations were tested and are shown in Table 2; for fleece material, both 15 cm and 23 cm lengths were used for the GEM-TRD_v2 and the Micromegas+GEM-TRD. Regarding Mylar foil, 15 cm of 30 m-thick foil with 240 m spacing was used on the GEM-TRD, while for the Micromegas+GEM-TRD 10 cm of 30 m-thick foil with 200 m spacing combined with 10 cm of 25 m-thick foil with 200 m spacing for a total length of 21 cm was used.
@@ 3.3 Data Selection
For in-beam measurements, several observables are recorded using fADC-125 electronics: pulse amplitude (corresponding to deposited energy) and drift time of clusters across the drift gap. Event-by-event analyses were performed to study the detector response for electrons and hadrons, enabling determination of e/ separation power. Figure 4 shows an example of raw waveforms collected for electrons in both the GEM-TRD and the Micromegas+GEM-TRD. For analysis of in-beam measurements, tight selections are applied to external PID detectors to assure sufficiently high sample purity of charged particle types. For each of the TRD prototypes tested, other detectors in the setup are used as an external track reference. Tight selections are then applied to achieve a clean sample of charged-track events and ensure maximum detection efficiency of relevant charged particle species. The detection efficiency for each prototype is calculated by reconstructing the track along one axis using external detectors and looking for a hit in the prototype that matches the external track projection into the prototype’s active area. For the signal-detection threshold used in the analysis, a typical threshold of 100 fADC units is used where 1 fADC unit corresponds to about 0.04 fC, or 250 electrons.